OpenStudy (anonymous):

a) write the equation of the median drawn to AE in point-slope form. b) write the equation of the perpendicular bisector drawn to AE in point slope form.

OpenStudy (anonymous):



OpenStudy (anonymous):

Have you attempted a or b yet?

OpenStudy (anonymous):

yes but i can't remember the equations

OpenStudy (anonymous):

i could really use the help

OpenStudy (ranga):

Find the midpoint of AE: ( (x1+x2)/2, (y1+y2)/2 ) Then find the equation of the median passing through C (-5,2) and the midpoint.

OpenStudy (anonymous):

The two point formula is:\[ \frac{y-y_1}{x-x_1} = \frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (anonymous):

You plug in values for everything except \(x\) and \(y\) and then isolate \(y\).

OpenStudy (anonymous):

for the midpoint i get (7,1)

OpenStudy (ranga):

Yes. Now find the equation of the line passing through (7,1) and (-5,2)

OpenStudy (anonymous):

\[y=-\frac{ 1 }{ 12 }x+\frac{ 19 }{ 12 }\]

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

then how do I get part b

OpenStudy (ranga):

Find the slope of AE. The slope of the line perpendicular to AE will be the negative reciprocal of the slope of AE. You already know the midpoint of AE (found in a) Knowing the slope of the perpendicular bisector and knowing the midpoint you can find the equation.